OpenStudy (anonymous):

Find cos θ if sin θ = -12/13and tan θ > 0.

OpenStudy (jdoe0001):

\(\bf sin(\theta) = -\cfrac{12}{13} \implies \cfrac{opposite}{hypotenuse} \implies \cfrac{b}{c}\\ c^2 = a^2 + b^2 \implies \sqrt{c^2-b^2} = a\) find "a" to get the cosine of the angle